# UFO: A Unified Approach to Fine-grained Visual Perception via Open-ended Language Interface

### v2

@@ 1 Introduction
Figure 1: Methods to augment MLLMs with fine-grained perception tasks. (a) Relying on task decoders [33, 68], (b) Previous text-based methods represent boxes with location tokens [46, 61] and represent masks with suboptimal polygons [66, 61] or textual classes [61, 34], (c) Ours: predicting open-ended text sequences while using a simple yet effective embedding retrieval approach to support masks.
@@
A straightforward approach adopted by many existing works [33, 68, 51, 54, 47] is to augment MLLMs with task decoders. As shown in Figure 1 (a), LISA [33] introduces SAM [30] to support segmentation. Similarly, VisionLLM v2 [68] employs extra box and mask decoders. However, this combination introduces several limitations. First, task decoders add architectural complexity, necessitating compatibility among multiple components whenever the LLM is scaled up or new task decoders are introduced. Second, it complicates the end-to-end training. For example, the last stage in VisionLLM v2 [68] is dedicated to finetuning the task decoders because they fail to converge in earlier stages. These issues create a significant discrepancy with traditional vision-language modeling, limiting their broader application in general-purpose MLLMs. To remove task decoders, another line of research [46, 6, 72, 66] adopt text-based methods, converting boxes into location tokens or textual numbers. However, many text-based methods do not support segmentation due to the challenges of encoding masks of arbitrary shapes and complexity [3] with limited text tokens.
@@
To address the above limitations, GiT [61] designs a language interface to support both box and mask outputs without task decoders. However, it still falls short of unifying fine-grained perception tasks into MLLMs. For segmentation, GiT utilizes two suboptimal representations, illustrated in Figure 1 (b). In instance segmentation, it predicts polygon coordinates, which introduce quantization errors. In semantic segmentation, it predicts textual classes for each pixel, resulting in overly long sequences. For detection, GiT translates boxes into four location tokens and restricts the output space exclusively to these tokens, incompatible with other open-ended vision-language tasks that generate text of arbitrary length. In addition, GiT does not scale to MLLM and uses a Vision Transformer (ViT [21]) for multimodal tasks, resulting in poor language abilities. Hence, it is essential to develop a more effective approach to unify fine-grained perception into MLLMs. This method should effortlessly align with open-ended language interfaces, involve minimal structural complexity and deliver excellent performance.
@@
In this paper, we present UFO, which unifies fine-grained perception tasks through the same open-ended language interface as vision-language tasks, without any task decoders. By translating all task outputs into open-ended text sequences, we demonstrate that competitive performance can be achieved without complex task-specific designs. As illustrated in Figure 1 (c), we reformulate segmentation as an embedding retrieval problem, where the mask token embedding computes similarity with image features by dot product, retrieving high-similarity positions to generate the mask. This design effectively leverages the output image features processed by MLLMs, which are often overlooked in previous methods. Our intuition is that since MLLMs achieve strong visual understanding, the mask information is already in the image features and we just need to retrieve it. Furthermore, we introduce a novel method that upsamples output masks by predicting multiple mask tokens, resulting in more refined masks and improved performance.
@@
We first validate our method following GiT [61], which uses a lightweight ViT for multimodal processing but can share the same formulation as MLLMs (see Table 1), allowing efficient validation. GiT constructs a comprehensive multi-task benchmark, which covers various granularity fine-grained perception tasks. Under the same evaluation protocols, UFO outperforms GiT by 12.3 mAP and 3.3 mIoU in COCO instance segmentation and ADE20K semantic segmentation, respectively. We then scale our method to larger MLLMs. Experimental results demonstrate that UFO effectively integrates fine-grained perception capabilities into MLLMs, achieving strong performance without task decoders. Leveraging the powerful language capabilities of MLLMs, UFO can extend to more language-dependent tasks, such as reasoning segmentation.
@@
(1) We introduce UFO, the first framework to unify diverse fine-grained perception tasks through the same open-ended language interface as vision-language tasks, without relying on task-specific decoders.
(2) We reformulate segmentation as an embedding retrieval problem, exploring both text generation and image representation abilities of the language interface, significantly outperforming previous text-based methods.
(3) Our framework seamlessly integrates with MLLMs, delivering competitive performance with a wide range of fine-grained perception tasks of varying difficulties.
@@ 2.1 Multimodal Large Language Models
Inspired by the success of large language models (LLMs), multimodal large language models (MLLMs) have rapidly advanced in recent years. Early efforts like LLaVA [39], MiniGPT-4 [81], and Instruct-BLIP [17] fine-tune LLMs with instruction-following datasets, demonstrating strong multimodal understanding. More advanced MLLMs like DeepSeek-VL [41], Qwen2-VL [65], and InternVL2 [12] have emerged recently, offering superior multimodal comprehension through larger model sizes and extensive training data. However, these models remain limited to image-level vision-language tasks, with less exploration of fine-grained visual perception, thereby restricting their application scope.
@@ 2.2 Extend MLLMs with Fine-grained Perception
Extend MLLMs with Task Decoders. Recent works [33, 68, 51, 54, 2, 78, 47, 69, 79] introduce task decoders to extend MLLMs with tasks like detection and segmentation. These models treat the MLLM as a coarse proposal extractor, passing the task-relevant embeddings to specialized decoders. The decoders then manage task-specific details, such as regressing boxes or generating masks. Although this approach yields strong performance, extra task decoders complicate architectures and training, undermining the unified design of MLLMs and limiting their potential.
Extend MLLMs with Text Outputs. To augment MLLMs with object-level perception tasks, previous methods [46, 6, 72, 60, 66, 5, 48] have employed location tokens or textual numbers to represent boxes. However, these approaches face challenges with pixel-level tasks such as segmentation, often requiring polygonal approximations [66, 48]. Although VistaLLM [48] reduces the errors of polygons through adaptive sampling and achieves competitive performance in referring segmentation, it remains inadequate for general segmentation tasks. First, polygons struggle to accurately represent “stuff” categories with amorphous regions (e.g., roads with parked cars), which are common in real world [80, 3, 74, 16]. Second, polygons inherently cause information loss, especially for detailed structures like retinal vessels [58]. In contrast, UFO leverages the multimodal outputs of MLLMs to generate precise masks for any shape, offering greater expressiveness and improved performance.
@@ 2.3 Vision Generalist Models
Vision generalist models aim to establish a unified framework supporting various vision-centric tasks. Inspired by the seq2seq framework in NLP, previous generalist models  [64, 61, 8, 42] transform visual tasks into sequence generation problems. Notably, GiT [61] unifies five core visual tasks through a language interface, supporting box, mask, and text outputs. However, these models typically focus solely on visual tasks and lack the advanced language capabilities required for complex multimodal tasks.
Figure 2: Overview of our approach. (a) Segmentation modeling: the mask token embedding retrieves similar image features to generate masks (shown with matching colors). (b) Upsampling masks by multiple mask tokens, retrieving more details by more tokens. We use $N$=2 to illustrate while using $N$=4 in implementation. (c) We output open-ended text sequences with textual numbers for detection.
@@
UFO-InternVL2-8B InternViT [12],MLP InternLM2.5 Tokenizer [77] InternLM2.5-7B [77]
Table 1: We abstract current multimodal architectures into three components: (1) Image tokenizer, converting images into visual tokens; (2) Text tokenizer, outputting text tokens; (3) Multimodal transformer, jointly processing visual and text tokens. We construct three variants by this formulation and list their architecture details.
@@ 3.1 Preliminary
Our goal is to unify fine-grained perception tasks into the open-ended language interface, thereby ensuring compatibility with any multimodal architecture that supports the same interface. We abstract existing multimodal architecture into three components based on the modalities they process: image tokenizer, text tokenizer and multimodal transformer, as shown in Tabel 1. For example, in LLaVA [39], the image tokenizer includes a vision encoder and MLP connector that extract visual features and map them into the LLM’s input space, while the multimodal transformer corresponds to the LLM. This abstraction applies not only to MLLMs with various image tokenizers [39, 35, 20] but also to vision generalist models with similar architectures [61], significantly broadening the scope of our method. To avoid confusion, we will refer to MLLMs by default in the following sections.
@@ 3.2 Bounding Box Representation
To align with the open-ended language interface while avoiding the addition of extra location tokens, we directly translate boxes into textual numbers. Each box is represented by the coordinates of its top-left $(x_{1},y_{1})$ and bottom-right $(x_{2},y_{2})$ corners. The continuous values of these coordinates are discretized into integers within [0, range], enclosed by <box> and </box> tokens. If a class label is required, we simply prepend the textual class before the <box> token. For example, a box of a person can be represented as:
person,<box>465,268,589,344</box>
This approach uses open-ended sequences to represent boxes, effectively aligning with vision-language tasks.
@@ 3.3 Mask Representation
Segmentation by Embedding Retrieval. To incorporate the segmentation task using only the language interface, we reformulate it as an embedding retrieval problem. Specifically, we approach segmentation via mask prediction. We first augment the basic vocabulary of the model with a <MASK> token, which serves as the indicator for mask generation. When performing segmentation, the model is trained to output the <MASK> token, as shown in Figure 2 (a). Formally, given an input image $\mathbf{x_{v}}$ and a segmentation prompt $\mathbf{x_{t}}$, the model $\mathcal{F}$ generates the text response $\mathbf{y_{t}}$ and corresponding output embeddings $\mathbf{h_{t}}$, image features $\mathbf{h_{v}}$ as:
@@
Our approach removes the need for task decoders while leveraging the inherent capabilities of the MLLMs to effectively support segmentation tasks. We hypothesize that, in well-encoded image features, features sharing the same semantics will group into clusters. Therefore, generating a mask token embedding equates to identifying the center of the relevant image feature cluster, while computing the similarity between embeddings reflects this relationship.
Upsampling by Multiple Mask Tokens. Due to the redundancy in visual information, it is common to process visual features at reduced resolutions. For example, the CLIP-L/14 [49] model downsamples image features by a factor of 14. In our segmentation method, similarities are computed using downsampled image features, resulting in low-resolution masks. However, directly upsampling by interpolation leads to coarse results and suboptimal performance due to the high interpolation factor.
@@
Our method effectively leverages mask token embeddings as upsampling parameters, offering greater flexibility than traditional methods like bilinear interpolation and transposed convolution. Bilinear interpolation uses non-learnable parameters, while transposed convolution allows for learnable parameters, the same parameters are applied to all images after training. In contrast, we use embeddings generated by the network as the parameters, which can be customized for each image. This approach enables the model to generate optimal upsampling parameters dynamically, enhancing flexibility while achieving better performance.
@@ 3.4 Multi-Task Data Template
Based on the above designs, we construct multi-task data templates for joint training. We classify tasks into two categories based on output prediction number: single-prediction tasks like visual grounding produce one box or mask, and multi-prediction tasks like object detection generate several boxes. Merging multiple outputs into a long sequence is inefficient and the order among them is hard to define, making autoregressive learning of the sequence difficult [7]. Inspired by GiT [61], we adopt a parallel decoding approach that splits multi-prediction tasks into independent subtasks, each handling one prediction in parallel. This strategy effectively accelerates inference and enhances task scalability.
Single prediction. For tasks only require a single prediction, our task template is:
<Text Prompt><Image><Text Response>
@@
Multiple predictions. To efficiently support multi-prediction tasks, we split complex tasks into independent subtasks with single prediction, enabling parallel decoding within a batch. The key to achieving parallelism is to ensure all subtasks are independent. Typically, multiple boxes and masks correspond to different locations. Therefore, we introduce local image features in the input to differentiate these sub-tasks, serving as visual prompts. The template is structured as follows:
<Text Prompt><Image><Local><Text Response>
where <Local> refers to local image features interpolated by grids sampled on the image. Specifically, we uniformly sample $M$ grid points on the whole image and interpolate the local image features at each grid location. We then match the $M$ grid points with $K$ predictions, assigning each prediction to the nearest grid point. As shown in Figure 3, example sub-sequences might be:
Detect cow, ...<Image><Local>cow, <box>... Segment cow, ...<Image><Local>cow, <MASK>...
The remaining $M-K$ grid points predict the end token. In this way, $M$ grid points correspond to $M$ independent subtasks. They share the text prompt and image features but have distinct local features and text responses, forming $M$ separate sub-sequences that can be predicted in parallel within a batch. By breaking down complex tasks into simple single-prediction subtasks, this approach enhances efficiency while simplifying learning. We adopt this template for object detection, instance segmentation and semantic segmentation.

### v3

@@ 1 Introduction
Figure 1: Methods to augment MLLMs with fine-grained perception tasks. (a) Relying on task decoders [37, 77], (b) Previous text-based methods represent boxes with location tokens [52] and represent masks with suboptimal polygons [74, 69] or textual classes [69, 38], (c) Ours: predicting open-ended text sequences while using a simple yet effective embedding retrieval approach for masks.
@@
Many existing works [37, 77, 57, 60, 53, 89, 11] augment MLLMs with task-specific decoders, such as LISA [37] using SAM for segmentation or VisionLLM v2 [77] adding box and mask decoders. However, this combination introduces several limitations. First, task decoders add architectural complexity, necessitating compatibility among multiple components whenever the LLM is scaled up or new task decoders are introduced. Second, it complicates the end-to-end training. For example, the last stage in VisionLLM v2 [77] is dedicated to finetuning the task decoders because they fail to converge in earlier stages. These issues create a significant discrepancy with traditional vision-language modeling, limiting their broader application in general-purpose MLLMs. To remove task decoders, another line of research [52, 7, 81, 74, 54] converts boxes into location tokens or textual numbers and transform masks into polygon vertices. However, using a limited number of vertices for masks introduces quantization errors, especially for masks with complex shapes and multiple regions.
@@
To address the above limitations, GiT [69] uses two mask representations: for instance segmentation, it uses polygons, and for semantic segmentation, it predicts textual classes for each pixel, as shown in Figure 1 (b). However, it still falls short of unifying fine-grained perception tasks into MLLMs. Firstly, although textual class can represent masks with any shape, it results in overly long sequences and slow inference. Secondly, to achieve better performance, GiT sets specific vocabularies for each task (e.g., detection can only output location tokens), which is incompatible with open-ended text generation. Finally, GiT does not scale to MLLMs and uses a Vision Transformer (ViT [21]) for multimodal tasks, resulting in poor language abilities. Hence, it is essential to develop a more effective approach to unify fine-grained perception into MLLMs. This method should effortlessly align with open-ended language interfaces, involve minimal structural complexity and deliver excellent performance.
@@
In this paper, we present UFO, which unifies fine-grained perception tasks through the same open-ended language interface as vision-language tasks, without any task decoders. By carefully organizing and translating all task outputs into open-ended text sequences, we demonstrate that competitive performance can be achieved without complex task-specific designs. As illustrated in Figure 1 (c), we reformulate segmentation as an embedding retrieval problem, where the mask token embedding computes similarity with image features by dot product, retrieving high-similarity positions to generate the mask. This design effectively leverages the output image features processed by MLLMs, which are often overlooked in previous methods. Our intuition is that since MLLMs achieve strong visual understanding, the mask information is already in the image features and we just need to retrieve it. Furthermore, we introduce a novel method that upsamples output masks by predicting multiple mask tokens, resulting in more refined masks and improved performance. Thanks to this strategy, we can efficiently and accurately represent masks of any shape using only 16 tokens.
@@
We first validate our method following GiT [69], which uses a lightweight ViT but can share the same formulation as MLLMs (see Table 1), allowing efficient validation. GiT constructs a comprehensive multi-task benchmark, which covers various granularity fine-grained perception tasks. Under the same evaluation protocols, UFO outperforms GiT by 12.3 mAP and 3.3 mIoU in COCO instance segmentation and ADE20K semantic segmentation (see Table 2). We then scale our method to MLLMs to integrate language abilities with fine-grained perception. As shown in Table 3, UFO achieves competitive results in visual grounding without decoders or polygon approximations. Benefiting from the shared representations of the open language interface, UFO can deeply unify textual reasoning and image segmentation, surpassing the previous state-of-the-art method on the challenging ReasonSeg [37] benchmarks by 6.2 gIoU (see Table 7).
@@
(1) We introduce UFO, a unified framework for diverse fine-grained perception tasks through the same open-ended language interface as vision-language tasks, without task-specific decoders.
(2) We reformulate segmentation as an embedding retrieval problem, exploring both text generation and image representation abilities of the language interface, significantly outperforming previous text-based methods on instance and semantic segmentation tasks.
(3) Our framework seamlessly integrates with MLLMs, delivering better performance than previous state-of-the-art methods on the ReasonSeg benchmarks.
@@ 2.1 Multimodal Large Language Models
Inspired by the success of large language models (LLMs), multimodal large language models (MLLMs) have rapidly advanced in recent years. Early efforts [43, 92, 17] finetune LLMs with instruction datasets, demonstrating strong multimodal understanding. More advanced MLLMs like Qwen2.5-VL [2], and InternVL2.5 [12] have emerged recently, offering superior multimodal comprehension through larger model sizes and extensive training data. However, these models mainly focus on image-level vision-language tasks, with less exploration of fine-grained visual perception.
@@ 2.2 Extend MLLMs with Fine-grained Perception
Extend MLLMs with Task Decoders. Recent works [37, 77, 57, 60, 3, 87, 53, 78, 88, 84, 89, 76] introduce task decoders to extend MLLMs with tasks like detection and segmentation. These models treat the MLLM as a coarse proposal extractor, passing the task-relevant embeddings to specialized decoders. The decoders then manage task-specific details, such as regressing boxes or generating masks. Although this approach yields strong performance, extra task decoders complicate architectures and training, undermining the unified design of MLLMs and limiting their potential. Recently, HiMTok [73] reformulates segmentation as mask image generation. However, this approach still requires training a specialized VQ decoder, increasing both training and structural complexity.
Extend MLLMs with Text Outputs. For object-level tasks, previous methods [52, 7, 81, 68, 74, 6, 54] have employed location tokens or textual numbers to represent boxes. For pixel-level tasks such as segmentation, a common format is polygonal approximation [74, 54]. Although VistaLLM [54] reduces the errors of polygons through adaptive sampling, it is inadequate for general segmentation tasks. First, polygons struggle to accurately represent “stuff" categories with amorphous regions (e.g., roads with parked cars), which are common in real world [90, 4, 16]. Second, polygons inherently cause information loss, especially for detailed structures like retinal vessels [64]. Text4Seg [38] directly predicts textual labels for image patches but still requires an additional refiner (e.g., SAM [34]) to achieve better performance. In contrast, UFO leverages the multimodal outputs of MLLMs to generate precise masks for any shape, offering greater expressiveness and improved performance.
@@ 2.3 Vision Generalist Models
Vision generalist models aim to establish a unified framework supporting various vision-centric tasks. Inspired by the seq2seq framework in NLP, previous generalist models  [72, 69, 9, 47] transform visual tasks into sequence generation problems. Notably, GiT [69] unifies five core visual tasks by language interface, supporting box, mask, and text outputs. However, these models typically focus solely on visual tasks and lack the advanced language capabilities required for complex reasoning [37].
Table 1: We abstract current multimodal architectures into three components: (1) Image tokenizer, converting images into visual tokens; (2) Text tokenizer, outputting text tokens; (3) Multimodal transformer, jointly processing visual and text tokens. We construct three variants by this formulation.
@@
UFO-InternVL2.5-8B InternViT [12],MLP InternLM2.5 Tokenizer [86] InternLM2.5-7B [86]
@@ 3.1 Preliminary
Our goal is to unify fine-grained perception tasks into the open-ended language interface, thereby ensuring compatibility with any multimodal architecture that supports the same interface. We abstract existing multimodal architecture into three components based on the modalities they process: image tokenizer, text tokenizer and multimodal transformer, as shown in Table 1. For example, in LLaVA [43], the image tokenizer includes a vision encoder and MLP connector that extract visual features and map them into the LLM’s input space, while the multimodal transformer corresponds to the LLM. This abstraction applies not only to MLLMs with various image tokenizers [43, 40, 20] but also to vision generalist models with similar architectures [69], significantly broadening the scope of our method. To avoid confusion, we will refer to MLLMs by default in the following sections.
@@ 3.2 Bounding Box Representation
To align with the open-ended language interface while avoiding the addition of extra location tokens, we directly translate boxes into textual numbers. Each box is represented by the coordinates of its top-left $(x_{1},y_{1})$ and bottom-right $(x_{2},y_{2})$ corners. The continuous values of these coordinates are discretized into integers within [0, range], enclosed by <box> and </box> tokens. If a class label is required, we simply prepend the textual class before the <box> token. For example, a box of a person can be represented as: person,<box>465,268,589,344</box>. This method converts boxes to open-ended sequences, effectively aligning with vision-language tasks.
Figure 2: Overview of our approach. (a) Segmentation modeling: the mask token embedding retrieves similar image features to generate masks (shown with matching colors). (b) Upsampling masks by multiple mask tokens, retrieving more details by more tokens. We use NN=2 to illustrate while using NN=4 in implementation. (c) We output open-ended text sequences with textual numbers for detection.
@@ 3.3 Mask Representation
Segmentation by Embedding Retrieval. To incorporate the segmentation task using only the language interface, we reformulate it as an embedding retrieval problem. We first augment the basic vocabulary of the model with a <MASK> token, which serves as the indicator for mask generation. When performing segmentation, the model is trained to output the <MASK> token, as shown in Figure 2 (a). Formally, given an input image $\mathbf{x_{v}}$ and a segmentation prompt $\mathbf{x_{t}}$, the model $\mathcal{F}$ generates the text response $\mathbf{y_{t}}$ and corresponding output embeddings $\mathbf{h_{t}}$, image features $\mathbf{h_{v}}$ as:
@@
Our approach leverages MLLMs’ inherent capabilities for segmentation without task decoders. We hypothesize that, in well-encoded image features, features with the same semantics will group into clusters. Therefore, generating a mask token embedding equates to identifying the center of the relevant image feature cluster, while computing the similarity reflects this relationship. This approach can easily apply to other pixel-level tasks, such as depth estimation (see Table 20 in the appendix).
Upsampling by Multiple Mask Tokens. Due to the redundancy in visual information, it is common to process visual features at reduced resolutions. For example, the CLIP-L/14 [56] downsamples image features by a factor of 14. In above method, similarities are computed using downsampled image features, resulting in low-resolution masks. However, directly upsampling by interpolation leads to coarse results and suboptimal performance due to the high interpolation factor.
@@
Our method effectively leverages mask token embeddings as upsampling parameters, offering greater flexibility than traditional methods like bilinear interpolation and transposed convolution. Bilinear interpolation uses non-learnable parameters, while transposed convolution allows for learnable parameters, the same parameters are applied to all images after training. In contrast, we use embeddings generated by the network as the parameters, which can be customized for each image. This approach enables the model to generate optimal upsampling parameters dynamically, enhancing flexibility while achieving better performance (see Table 7).
Note that our method is fully compatible with open-ended language interfaces. We refer “open-ended" as the capability to generate variable-length text sequences terminated by an end-of-sequence token. In our approach, while we require fixed-length <MASK> tokens for segmentation, the text generation itself remains open-ended. Only after the generation is complete, we check if the output contains <MASK> segments of the required length. Furthermore, the interaction with image features is also performed after text generation is finished.
@@ 3.4 Multi-Task Data Template
Based on the above designs, we construct multi-task data templates for joint training. We classify tasks into two categories based on prediction number: single-prediction tasks like grounding produce one box or mask, and multi-prediction tasks like object detection generate several boxes. Merging multiple outputs into a long sequence is inefficient and the order among them is hard to define, making autoregressive learning of the sequence difficult [8]. Therefore, we adopt a parallel decoding approach that splits multi-prediction tasks into independent subtasks, each handling one prediction in parallel. This strategy effectively accelerates inference and enhances task scalability.
Single prediction. For tasks only require a single prediction, our task template is: <Text Prompt><Image><Text Response>.
@@
Multiple predictions. To efficiently support multi-prediction tasks, we split complex tasks into independent subtasks with single prediction, enabling parallel decoding within a batch. The key to achieving parallelism is to ensure all subtasks are independent. Typically, multiple boxes and masks correspond to different locations. Therefore, we introduce local image features in the input to differentiate these sub-tasks, serving as visual prompts. The template is structured as follows: <Text Prompt><Image><Local><Text Response>, where <Local> refers to local image features interpolated by grids sampled on the image. The core idea is that each grid point is responsible for detecting its spatially nearest objects. During training, each ground-truth object is assigned to its nearest grid point, while the remaining grid points are assigned to predict end-of-sequence tokens. During inference, as illustrated in Figure 3, we sample points over the image, typically of size $K\times K$, resulting in a total of $M=K^{2}$ points. We then interpolate the image features at each of the $M$ grid locations to extract distinct grid features. These grid features, along with the global image features and the text prompt, are fed into the LLM. The example input sequence is structured as follows:
$\text{Detect cow, person, duck.}<\mathrm{Image}><\mathrm{Local}_{1}><\mathrm{Local}_{2}>\ldots<\mathrm{Local}_{M}>$ (6)
To enforce the independence of predictions for each grid point, we modify the self-attention mask to isolate each grid feature from the others. This ensures that the generation for one point does not influence another. Then we start generating in an autoregressive manner, with the difference that we predict M tokens at each forward step instead of just one. The generation process looks like this:
$\mathrm{Local}_{1}>\ldots<\mathrm{Local}_{M}>|<T^{1}_{1}>\ldots<T^{1}_{M}>|<T^{2}_{1}>\ldots<T^{2}_{M}>|\ldots$ (7)
where $T^{j}_{i}$ denotes the $j$-th generated token for the $i$-th grid sequence, and | distinguishes the tokens produced in each forward pass. This decoding strategy shares the same philosophy as blockwise prediction [65] in LLMs, accelerating inference by generating multiple tokens simultaneously.
After decoding, we obtain $M$ output sequences. For detection, some might identify objects (Duck, <box>... or Cow, <box>...), while others corresponding to empty regions will predict end-of-sequence tokens. For instance segmentation, the process is identical, with the textual box representations (<box>...) being replaced by mask tokens (<MASK>...). This approach not only enhances efficiency but also simplifies the problem by breaking it down into simple, localized prediction tasks.